OpenStudy (anonymous):

Solve for x 9.1x - 5 = 8.1x I've figured out that x is -5, but when I check my work the answer it's wrong: 9.1(-5) - 5 = 8.1(-5) = -45.5 - 5 = -40.5 I'm so confused.

OpenStudy (anonymous):

no, the answer is x = 5. check it.

OpenStudy (anonymous):

I kinda understand, but if you could explain it to me I'd appreciate it.

OpenStudy (anonymous):

simple, 9.1x-5=8.1x 9.1x-8.1x=5 then use checking :)

OpenStudy (anonymous):

u probably didn't transpose it?

OpenStudy (anonymous):

I get it: -45.5 + 5 = 40.5

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

a million thank yous

OpenStudy (anonymous):

no problem :)